# D-Gril: End-to-End Topological Learning with $2$-parameter Persistence

### v2

@@ 1 Introduction
In recent years, persistent homology, one of the flagship concepts of Topological Data Analysis (TDA), has found its use in many fields such as neuroscience, material science, sensor networks, shape recognition, gene expression data analysis and many more Giunti et al. (2022). The performance of machine learning models such as Graph Neural Networks (GNNs) can be enhanced by augmenting topological information captured by persistent homology (Hofer et al., 2017; Dehmamy et al., 2019; Carrière et al., 2020; Horn et al., 2022). Classical persistent homology, also known as $1$-parameter persistence, captures the evolution of topological structures in a simplicial complex $K$ following a filter function $f\colon K\to\mathbb{R}$. The evolution of topological structures, in this case, can be completely characterized and compactly represented as persistence diagrams or equivalently barcodes Zomorodian & Carlsson (2004); Edelsbrunner et al. (2002). These persistence diagrams or barcodes can be vectorized Bubenik (2015); Reininghaus et al. (2015); Adams et al. (2017); Hofer et al. (2019); Carrière et al. (2020); Kim et al. (2020) and used in machine learning pipelines. In most applications, the simplicial complex $K$ is given and the choice of the filter function $f$ depends on the application. Choosing an appropriate filter function can be challenging. To avoid this, in Hofer et al. (2020), the authors proposed an end-to-end learning framework to learn the filter function rather than relying on making the right choice. They showed that learning the filter function performs better than the standard choices of filter functions on many graph datasets.
@@
To the best of our knowledge, end-to-end topological learning with multiparameter persistence has not been investigated in the literature. In concurrent work, Scoccola et al. (2024) address this question by developing a general framework suitable for a class of invariants of multiparameter persistence. Our approach is specific to GRIL but is more direct and yields a simpler line of argument.
@@ 3.1 Multiparameter persistent homology
We refer the reader to MacLane (1971) for the definitions of limit, colimit, and the construction of the canonical limit-to-colimit map. Intuitively, generalized rank captures the number of independent topological features supported over the interval $I$. It can be computed efficiently for $2$-parameter persistence modules using an algorithm in Dey et al. (2024).
@@ 3.2 Stochastic sub-gradient descent
We note that all semi-algebraic functions are definable. In fact, the author in, Wilkie (1996), shows that there exists an o-minimal structure that simultaneously contains all semi-algebraic sets and the graph of the exponential function. The authors in Davis et al. (2020) use this result to show that a neural network, being a composition of definable functions, is definable (Corollary 5.11 Davis et al. (2020)). For readers not familiar with the notion of definable functions, it is sufficient to know that a piecewise affine map is definable as it is semi-algebraic.
@@ 4 Differentiability of Gril
We begin this section by proving that Gril is a piecewise affine map in section 4.1. Gril being piecewise affine leads us to the fact that stochastic sub-gradient converges almost surely on Gril, which is discussed in section 4.2. In section 4.3, we give an explicit formula for the differential of Gril. We conclude this section by pointing out some practical issues that arise while implementing this differentiable pipeline and discussing the solutions in section 4.4.
@@ 4.1 Gril as a piecewise affine map
By Rademacher’s theorem Federer (2014), the above proposition also says that $\Lambda^{\mathbf{p}}_{k,\ell}$ is differentiable almost everywhere.
@@
Corresponding to each such pair of simplex coordinates, we get one hyperplane. Here, in the example, we chose one simplex coordinate to be $x$-coordinate and the other to be $y$-coordinate. This, however, need not always be the case: we can have all four possible combinations of $x$- and $y$-coordinates, corresponding to each of which we get a hyperplane. The exact formula of all such hyperplanes is given in Appendix A. Combining all these hyperplanes, we get an arrangement $\mathcal{H}$ of hyperplanes in $\mathbb{R}^{2n}$ Goodman & O’Rourke (2004). The arrangement $\mathcal{H}$ partitions $\mathbb{R}^{2n}$ into relatively open 222i.e., open relative to their topological closure in $\mathbb{R}^{2n}$. $r$-cells, $r\in\{0,\ldots,2n\}$. We observe that this arrangement induces an affine stratification $\mathcal{S}_{\mathcal{H}}$ (refer Leygonie et al. (2021) for the formal definition) of $\mathbb{R}^{2n}$ where the $r$-dimensional strata are precisely the $r$-cells.
@@
In the following we prove that $\mathcal{G}$ is a piecewise affine map relative to this arrangement, meaning that it is affine on each stratum of $\mathcal{S}_{\mathcal{H}}$. To this end, we introduce the notions of upper boundary, lower boundary, and constraining simplex coordinate for an $\ell$-worm. These notions will allow us to characterize the strata of $\mathcal{S}_{\mathcal{H}}$ in terms of conditions on the bifiltration function. We give an intuitive explanation of these concepts below using Figure 1. Formal definitions are given in Appendix A.
In Figure 1, a $2$-worm is shown with its lower boundary colored in blue and its upper boundary in red. As depicted, the lower boundary consists of the part of the boundary that is in the lower half of the worm, while the upper boundary consists of the rest of the boundary. More precisely, a point belongs to the lower boundary (resp. upper boundary) if the open lower-set (resp. upper-set) of the point does not intersect the worm.
@@ 4.2 Stochastic sub-gradient descent
Figure 2: An intuitive understanding of the gradient assignment described in Theorem 4.7. The $y$-coordinate $\sigma^{y}$ is at a distance of $d$ from the $y$-coordinate of $\mathbf{p}$ in the left figure; $\sigma$ is the only constraining simplex for the $2$-worm. In the figure on the right, $\sigma^{y}$ has reduced by $\epsilon$ and is now at a distance $d+\epsilon$ from $\mathbf{p}^{y}$. As a consequence, the value of Gril increases from $d$ to $d+\epsilon$. Thus, $\frac{\partial\Lambda^{\mathbf{p}}_{k,\ell}(\mathbf{v}_{f})}{\partial\sigma^{y}}=-1$ as described in Theorem 4.7.
@@ 4.3 Differential of Gril
$\frac{\partial\Lambda^{\mathbf{p}_{j}}_{k,\ell}(\mathbf{v}_{f})}{\partial\mathbf{p}_{j}^{x}}=\begin{cases}+1,&\text{if there exists a lower x-constraining simplex for }\boxed{\mathbf{p}_{j}}^{\ell}_{d_{j}},\\ -1,&\text{if there exists an upper x-constraining simplex for }\boxed{\mathbf{p}_{j}}^{\ell}_{d_{j}},\\ 0,&\text{otherwise,}\end{cases}$
$\frac{\partial\Lambda^{\mathbf{p}_{j}}_{k,\ell}(\mathbf{v}_{f})}{\partial\mathbf{p}_{j}^{y}}=\begin{cases}+1,&\text{if there exists a lower y-constraining simplex for }\boxed{\mathbf{p}_{j}}^{\ell}_{d_{j}},\\ -1,&\text{if there exists an upper y-constraining simplex for }\boxed{\mathbf{p}_{j}}^{\ell}_{d_{j}},\\ 0,&\text{otherwise}.\end{cases}$
@@ 5.1 Experimental Setup
Every instance in each dataset is an attributed graph with a label associated with it. We use a Graph Isomorphism Network (GIN) Xu et al. (2019) to obtain a bifiltration function over the vertices. We consider the lower-star bifiltration of this function. This gives us a bifiltration function on the entire graph, which we denote as $f$. Let $\mathbf{v}_{f}$ denote the vector representation of $f$ which we get after ordering the simplices of the graph. We divide the range of $f$ into a grid of size $100\times 100$. We randomly initialize the $s$ center points from this grid and compute the vector $\mathcal{G}_{k,\ell}(\mathbf{v}_{f})$ (Eq. (2)) for $k=1,2,3$ and $\ell=2$ at these center points. For computational efficiency, we restrict the center points and the boundaries of the worms to the grid. Note that the bifiltration function values are not restricted to the grid. This ensures that the differential of Gril is computed in $\mathbb{R}^{2n}$ and the steps of stochastic sub-gradient descent are updated in $\mathbb{R}^{2n}$. This vector $\mathcal{G}_{k,\ell}(\mathbf{v}_{f})$ is fed to a $3$-layer Multilayer Perceptron (MLP) classifier and the loss is back-propagated through the Gril computation according to the gradient described in Section 4. Note that the positions of the $s$ center points are also optimized as the gradient can be backpropagated to the coordinates the center points as well. Additional details and hyperparameter choices are given in Appendix B.
@@ 5.2 Bio-activity Prediction Datasets
The data is extracted from the ChEMBL database Gaulton et al. (2012); Davies et al. (2015). Each dataset contains the SMILES Weininger (1988) encoding of a novel drug (compound), and activity pairs for a target of interest. For example, the EGFR dataset contains all the molecules that have been tested against epidermal growth factor receptor (EGFR) kinase, and their measured bio-activity. The SMILES encodings of the molecules are then converted to graphs with Molfeat Noutahi et al. (2023). Bio-activity is measured in half maximal inhibitory concentration $(IC_{50})$, which qualitatively indicates how much a drug is needed, in vitro, to inhibit a particular process by $50\%$. To facilitate the comparison of $IC_{50}$ values, it is common practice to convert $IC_{50}$ to $pIC_{50}=-\log_{10}(IC_{50})$, expressed in molar units. Threshold for activity cutoff $pIC_{50}=6.3$ is used throughout the paper. More details about the datasets are given in Appendix B.
Dataset GIN GIN-Gril D-Gril
EGFR $55.60\pm 8.61$ $58.39\pm 2.51$ $\mathbf{65.38\pm 5.34}$
ERRB2 $57.06\pm 8.58$ $61.28\pm 3.66$ $\mathbf{69.24\pm 5.06}$
CHEMBL1163125 $58.48\pm 4.37$ $54.13\pm 1.09$ $\mathbf{65.69\pm 1.57}$
CHEMBL2148 $52.88\pm 0.91$ $50.24\pm 0.18$ $\mathbf{53.13\pm 3.70}$
CHEMBL4005 $55.90\pm 4.63$ $51.85\pm 3.55$ $\mathbf{59.34\pm 5.70}$
@@
In the first set of experiments, we compare D-Gril with (i) a standard GNN model–GIN with sum-pooling, (ii) with GIN and Gril as a readout layer. Note that for the results on GIN-Gril, Gril is used as a passive readout layer, i.e., we obtain a bifiltration function from a pre-trained GIN (pre-trained for graph classification on the same dataset) and compute Gril on it. We use the exact same pipeline as GIN-Gril and train it end-to-end using D-Gril and report the results in the last column. The classifier, $3$-layer MLP, is the same for all the experiments. Refer to Figure 3 for the pipeline. From Table 1, we can see that adding topological information after GIN has been trained can improve the performance in some cases and need not be beneficial in others. However, it is clear from the table that adding topological information in an end-to-end framework appears to be beneficial for bio-activity prediction.
Further, we perform a series of experiments where we augment other popular molecular fingerprints such as ECFP and Morgan3 (Rogers & Hahn, 2010; Morgan, 1965) with the D-Gril framework. In these experiments, we compare the performance of the model with and without D-Gril augmentation. We report the results in Table 2. We can see from the table that augmenting with D-Gril seems to, generally, increase the performance of the model. Indeed, fingerprints such as ECFP and Morgan3 can essentially represent an infinite number of different molecular features, including stereochemical information (Rogers & Hahn, 2010) but they are not effective in capturing global features of molecules such as size and shape (Capecchi et al., 2020). Combining D-Gril with these fingerprints augments the model with topological information, leading to an improved performance.
@@
Dataset D-Gril Gril MP-I MP-L MP-K P
MUTAG $\mathbf{85.09\pm 5.99}$ $83.49\pm 3.64$ $74.99\pm 2.79$ $82.42\pm 3.72$ $79.27\pm 2.45$ $66.50\pm 0.87$
PROTEINS $69.45\pm 4.11$ $66.31\pm 2.34$ $70.80\pm 3.09$ $\mathbf{70.80\pm 1.31}$ $61.70\pm 2.98$ $59.57\pm 0.08$
DHFR $61.24\pm 4.37$ $\mathbf{61.64\pm 1.66}$ $60.98\pm 0.10$ $61.11\pm 0.25$ $60.98\pm 0.10$ $60.98\pm 0.10$
COX2 $78.16\pm 0.41$ $78.16\pm 0.41$ $78.16\pm 0.41$ $78.16\pm 0.41$ $78.16\pm 0.41$ $78.16\pm 0.41$
IMDB-BINARY $\mathbf{58.90\pm 1.90}$ $50.00\pm 0.00$ $56.60\pm 2.94$ $50.00\pm 0.00$ $50.30\pm 1.12$ $50.00\pm 0.00$

### v3

@@ 1 Introduction
In recent years, persistent homology, one of the flagship concepts of Topological Data Analysis (TDA), has found its use in many fields such as neuroscience, material science, sensor networks, shape recognition, gene expression data analysis and many more Giunti et al. (2022); Coskunuzer & Akçora (2024). The performance of machine learning models such as Graph Neural Networks (GNNs) can be enhanced by augmenting topological information captured by persistent homology (Hofer et al., 2017; Dehmamy et al., 2019; Carrière et al., 2020; Horn et al., 2022). Classical persistent homology, also known as $1$-parameter persistence, captures the evolution of topological structures in a simplicial complex $K$ following a filter function $f\colon K\to\mathbb{R}$. The evolution of topological structures, in this case, can be completely characterized and compactly represented as persistence diagrams or equivalently barcodes Zomorodian & Carlsson (2004); Edelsbrunner et al. (2002). These persistence diagrams or barcodes can be vectorized Bubenik (2015); Reininghaus et al. (2015); Adams et al. (2017); Hofer et al. (2019); Carrière et al. (2020); Kim et al. (2020) and used in machine learning pipelines. In most applications, the simplicial complex $K$ is given and the choice of the filter function $f$ depends on the application. Choosing an appropriate filter function can be challenging. To avoid this, in Hofer et al. (2020), the authors proposed an end-to-end learning framework to learn the filter function rather than relying on making the right choice. They showed that learning the filter function performs better than the standard choices of filter functions on many graph datasets.
@@
In a recent work, Scoccola et al. (2024) present an end-to-end topological learning method by developing a general framework suitable for a class of invariants of multiparameter persistence. Our approach though specific to Gril is more direct, uses simpler arguments, and is faster for training.
@@ 3.1 Multiparameter persistent homology
We refer the reader to MacLane (1971) for the definitions of limit, colimit, and the construction of the canonical limit-to-colimit map. Intuitively, generalized rank captures the number of independent topological features supported over the interval $I$. It can be computed efficiently for $2$-parameter persistence modules Dey et al. (2024) using zigzag persistence Dey & Hou (2022).
@@ 3.2 Stochastic sub-gradient descent
We note that all semi-algebraic functions are definable. In fact, the author in, Wilkie (1996), shows that there exists an o-minimal structure that simultaneously contains all semi-algebraic sets and the graph of the exponential function. The authors in Davis et al. (2020) use this result to show that a neural network, being a composition of definable functions, is definable (Corollary 5.11 Davis et al. (2020)). For readers not familiar with the notion of definable functions, it is sufficient to know that a piecewise affine map is definable.
@@ 4 Differentiability of Gril
We begin this section by proving that Gril is a piecewise affine map in section 4.1. Gril being piecewise affine leads us to the fact that stochastic sub-gradient converges almost surely on Gril, which is discussed in section 4.2. In section 4.3, we give an explicit formula for the differential of Gril.
@@ 4.1 Gril as a piecewise affine map
By Rademacher’s theorem Federer (2014), $\Lambda^{\mathbf{p}}_{k,\ell}$ is differentiable almost everywhere.
@@
Corresponding to each such pair of simplex coordinates, we get one hyperplane. Here, in the example, we chose one simplex coordinate to be $x$-coordinate and the other to be $y$-coordinate. This, however, need not always be the case: we can have all four possible combinations of $x$- and $y$-coordinates, corresponding to each of which we get a hyperplane. The exact formula of all such hyperplanes is given in Appendix A. Combining all these hyperplanes, we get an arrangement $\mathcal{H}$ of hyperplanes in $\mathbb{R}^{2n}$ Goodman & O’Rourke (2004). The arrangement $\mathcal{H}$ partitions $\mathbb{R}^{2n}$ into relatively open $r$-cells, $r\in\{0,\ldots,2n\}$. We observe that this arrangement induces an affine stratification $\mathcal{S}_{\mathcal{H}}$ (refer Leygonie et al. (2021) for the formal definition) of $\mathbb{R}^{2n}$ where the $r$-dimensional strata are precisely the $r$-cells.
@@
In the following we prove that $\mathcal{G}$ is a piecewise affine map relative to this arrangement, meaning that it is affine on each stratum of $\mathcal{S}_{\mathcal{H}}$. To this end, we introduce the notions of upper boundary, lower boundary, and constraining simplex coordinate for an $\ell$-worm which will allow us to characterize the strata of $\mathcal{S}_{\mathcal{H}}$ in terms of conditions on the bifiltration function. We give an intuitive explanation of these concepts using Figure 1. Formal definitions are given in Appendix A.
In Figure 1, a $2$-worm is shown with its lower boundary colored in blue and its upper boundary in red. As depicted, the lower boundary consists of the part of the boundary that is in the lower half of the worm, while the upper boundary consists of the rest of the boundary, i.e., a point belongs to the lower boundary (resp. upper boundary) if the open lower-set (resp. upper-set) of the point does not intersect the worm.
@@ 4.2 Stochastic sub-gradient descent
Figure 2: An intuitive understanding of the gradient assignment described in Theorem 4.7. The $y$-coordinate $\sigma^{y}$ is at a distance of $d$ from the $y$-coordinate of $\mathbf{p}$ in the left figure; $\sigma$ is the only constraining simplex for the $2$-worm. In the figure on the right, $\sigma^{y}$ has reduced by $\epsilon$ and is now at a distance $d+\epsilon$ from $\mathbf{p}^{y}$. As a consequence, the value of Gril increases from $d$ to $d+\epsilon$. Thus, $\frac{\partial\Lambda^{\mathbf{p}}_{k,\ell}(\mathbf{v}_{f})}{\partial\sigma^{y}}=-1$.
@@ 4.3 Differential of Gril
$\frac{\partial\Lambda^{\mathbf{p}_{j}}_{k,\ell}(\mathbf{v}_{f})}{\partial\mathbf{p}_{j}^{x}}=\begin{cases}+1,&\text{if there exists a lower x-constraining}\\ &\text{simplex for }\boxed{\mathbf{p}_{j}}^{\ell}_{d_{j}},\\ -1,&\text{if there exists an upper x-constraining}\\ &\text{simplex for }\boxed{\mathbf{p}_{j}}^{\ell}_{d_{j}},\\ 0,&\text{otherwise,}\end{cases}$
$\frac{\partial\Lambda^{\mathbf{p}_{j}}_{k,\ell}(\mathbf{v}_{f})}{\partial\mathbf{p}_{j}^{y}}=\begin{cases}+1,&\text{if there exists a lower y-constraining}\\ &\text{simplex for }\boxed{\mathbf{p}_{j}}^{\ell}_{d_{j}},\\ -1,&\text{if there exists an upper y-constraining}\\ &\text{simplex for }\boxed{\mathbf{p}_{j}}^{\ell}_{d_{j}},\\ 0,&\text{otherwise.}\end{cases}$
@@ 4.4 Practical Considerations
We observe that if the bifiltration $f$ is extended in the manner described above to $\bar{f}$, then we end up in a scenario where multiple simplices have the same $x$ or $y$-values. This is a degenerate scenario with respect to the assumptions in Theorem 4.5. To avoid this, in practice, we add an infinitesimal perturbation to the higher dimensional simplices to ensure the values are not exactly the same. We note that this is, in spirit, similar to gradient sampling.
@@ 5.1 Experimental Setup
Every instance in each dataset is an attributed graph with a label associated with it. We use a Graph Isomorphism Network (GIN) Xu et al. (2019) to obtain a bifiltration function over the vertices. We consider the lower-star bifiltration of this function. This gives us a bifiltration function on the entire graph, which we denote as $f$. Let $\mathbf{v}_{f}$ denote the vector representation of $f$ which we get after ordering the $n$ simplices of the graph. We divide the range of $f$ into a grid of size $100\times 100$. We randomly initialize the $s$ center points from this grid and compute the vector $\mathcal{G}_{k,\ell}(\mathbf{v}_{f})$ (Eq. (2)) for $k=1,2,3$ and $\ell=2$ at these center points. This vector $\mathcal{G}_{k,\ell}(\mathbf{v}_{f})$ is fed to a $3$-layer Multilayer Perceptron (MLP) classifier and the loss is back-propagated through the Gril computation according to the gradient described in Section 4. Note that the positions of the $s$ center points are also optimized as the gradient can be backpropagated to the coordinates of the center points as well. Additional details can be found in Appendix B.
@@ 5.2 Bio-activity Prediction Datasets
The data is extracted from the ChEMBL database Gaulton et al. (2012); Davies et al. (2015). Each dataset contains the SMILES Weininger (1988) encoding of a novel drug (compound), and activity pairs for a target of interest. For example, the EGFR dataset contains all the molecules that have been tested against epidermal growth factor receptor (EGFR) kinase, and their measured bio-activity. The SMILES encodings of the molecules are then converted to graphs with Molfeat Noutahi et al. (2023). Bio-activity is measured in half maximal inhibitory concentration $(IC_{50})$, which qualitatively indicates how much a drug is needed, in vitro, to inhibit a particular process by $50\%$. To facilitate the comparison of $IC_{50}$ values, it is common practice to convert $IC_{50}$ to $pIC_{50}=-\log_{10}(IC_{50})$, expressed in molar units. Threshold for activity cutoff $pIC_{50}=6.3$ is used throughout the paper. Additional details are given in Appendix B.
Dataset GIN GIN-Gril GIN-Mpsm D-Gril
EGFR $55.60\pm 8.61$ $58.39\pm 2.51$ $\mathbf{66.97\pm 1.87}$ $65.38\pm 5.34$
ERRB2 $57.06\pm 8.58$ $61.28\pm 3.66$ $68.75\pm 1.98$ $\mathbf{69.24\pm 5.06}$
 CHEMBL1163125 $58.48\pm 4.37$ $54.13\pm 1.09$ $61.23\pm 2.47$ $\mathbf{65.69\pm 1.57}$
CHEMBL2148 $52.88\pm 0.91$ $50.24\pm 0.18$ $51.45\pm 0.87$ $\mathbf{53.13\pm 3.70}$
CHEMBL4005 $55.90\pm 4.63$ $51.85\pm 3.55$ $\mathbf{61.35\pm 1.49}$ $59.34\pm 5.70$
@@
In the first set of experiments, we compare D-Gril with (i) a standard GNN model–GIN with sum-pooling, (ii) with GIN and Gril as a readout layer and with (iii) GIN and Mpsm Scoccola et al. (2024). Note that for the results on GIN-Gril, Gril is used as a passive readout layer, i.e., we obtain a bifiltration function from a pre-trained GIN (pre-trained for graph classification on the same dataset) and compute Gril on it. We use the exact same pipeline as GIN-Gril and train it end-to-end using D-Gril and GIN-Mpsm and report the results. The classifier, $3$-layer MLP, is the same for all the experiments. Refer to Figure 3 for the pipeline. From Table 1, we can see that adding topological information after GIN has been trained can improve the performance in some cases and need not be beneficial in others. However, it is clear that adding topological information in an end-to-end framework appears to be beneficial for bio-activity prediction. The performance of D-Gril is comparable to Mpsm.
Further, we perform a series of experiments where we augment other popular molecular fingerprints such as ECFP and Morgan3 (Rogers & Hahn, 2010; Morgan, 1965) with the D-Gril framework. In these experiments, we compare the performance of the model with and without D-Gril augmentation. We report the results in Table 2. We can see from the table that augmenting with D-Gril seems to, generally, increase the performance of the model. Fingerprints such as ECFP and Morgan3 can essentially represent an infinite number of different molecular features, including stereochemical information (Rogers & Hahn, 2010) but they are not effective in capturing global features of molecules such as size and shape (Capecchi et al., 2020). Combining D-Gril with these fingerprints augments the model with topological information, yielding an improved performance.
@@
Dataset D-Gril Mpsm Gril MP-I MP-L MP-K P
MUTAG $\mathbf{85.09\pm 5.99}$ $78.44\pm 3.33$ $83.49\pm 3.64$ $74.99\pm 2.79$ $82.42\pm 3.72$ $79.27\pm 2.45$ $66.50\pm 0.87$
PROTEINS $69.45\pm 4.11$ $68.37\pm 3.31$ $66.31\pm 2.34$ $70.80\pm 3.09$ $\mathbf{70.80\pm 1.31}$ $61.70\pm 2.98$ $59.57\pm 0.08$
DHFR $61.24\pm 4.37$ $60.97\pm 1.98$ $\mathbf{61.64\pm 1.66}$ $60.98\pm 0.10$ $61.11\pm 0.25$ $60.98\pm 0.10$ $60.98\pm 0.10$
COX2 $78.16\pm 0.41$ $77.73\pm 0.36$ $78.16\pm 0.41$ $78.16\pm 0.41$ $78.16\pm 0.41$ $78.16\pm 0.41$ $78.16\pm 0.41$
IMDB-BINARY $\mathbf{62.60\pm 6.56}$ $60.80\pm 3.82$ $50.00\pm 0.00$ $56.60\pm 2.94$ $50.00\pm 0.00$ $50.30\pm 1.12$ $50.00\pm 0.00$
@@
Dataset D-Gril Mpsm
MUTAG 00:05:08 00:08:47
COX2 00:11:34 00:18:32
DHFR 00:16:48 00:27:12
PROTEINS 00:30:03 00:50:26
Table 4: Reported times (hh:mm:ss) are training times per fold averaged over 5 folds that we used for training. All the experiments have been performed on a machine with AMD EPYC 7313 16-Core Processor and NVIDIA A10 GPU.